OpenStudy (saphiraisbae123):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y A. yes; k = 3/5 B. no C. yes; k = 5/3 D. yes; k = -5/3

OpenStudy (danjs):

if there is a direct variation for y and x, you can say y is proportional to x.

OpenStudy (danjs):

y is proportional and varies directly with x \[\large y~~\alpha~~x\] put in a proportionality constant k to make an equation. \[\large y=k*x\]

OpenStudy (danjs):

Solving this given equation for y 5x = -3y \[y=-\frac{ 5 }{ 3 }*x\] You see it is a direct variation y=k*x, with k being -5/3

OpenStudy (danjs):

as opposed to an inverse variation, if y varies inversely with x, then you have \[\large y~~\alpha~~\frac{ 1 }{ x }\] \[\large y=\frac{ k }{ x }\]